OpenStudy (anonymous):

I'm a bit confused... can someone help explain this to me? I will medal and fan (if I'm already a fan I will still medal). Write the quadratic variation equation. y varies directly with x^2 and y=15 when x=3

OpenStudy (agl202):

y = kx^2, where k is any constant y=15; x=3 15 = k(3)^2 15=9k k=15/9 y= 15x^2 / 9 y= 5x^2 / 3 Hope this helps! =)

OpenStudy (anonymous):

y = k(x^2) 9 = 6^2k 9 = 36 k k = 9/36 or 1/4 So, the equation is y = (1/4)x^2

OpenStudy (anonymous):

Thank-you Both! @kittenlover731 & @Agl202

OpenStudy (agl202):

Anytime : )